OpenStudy (anonymous):

The 12:00 p.m. eastbound train left the station at a constant speed of 40 miles per hour. How can I find the distance travelled? I know that the formula is speed x time but I can't apply it here because of the time. Please help.

OpenStudy (johnweldon1993):

You cannot solve it....you would need a time and you are correct in stating that d = speed * time with only speed...we have 2 unknowns...and we cannot solve for distance all you can do is make a generalized statement and say the distance traveled is = 40 * how ever many hours they travel d = 40t

OpenStudy (anonymous):

This was my word problem: The 12:00 p.m. eastbound train left the station at a constant speed of 40 miles per hour. At 12:45 p.m., the next eastbound train left the station at a constant speed of 60 miles per hour. Assuming neither train stops along the way, how far apart will the two trains be at 2:00 p.m.?

OpenStudy (anonymous):

I found the time for the second train which was 80 miles.

OpenStudy (anonymous):

How do I find the distance for the first train then? =)

OpenStudy (whpalmer4):

Okay, so you do know the time! The first train travels for 2 hours at 40 miles per hour, right?

OpenStudy (anonymous):

oh crap. thank you! i was lost for some reason.

OpenStudy (anonymous):

my bad!

OpenStudy (whpalmer4):

The second train travels at 60 miles per hour. It starts at 12:45 PM and goes until 2 PM. How much time elapses? Multiply that by 60 miles per hour...

OpenStudy (anonymous):

is the answer 10 miles?

OpenStudy (whpalmer4):

Well, how far does each train travel?

OpenStudy (whpalmer4):

First train goes 40 miles/hr * 2 hr = 80 miles Second train goes 60 miles/hr * 1 1/4 hr = ?

OpenStudy (anonymous):

my bad again! it's 80 mph because it'll be speed x time

OpenStudy (whpalmer4):

no, not 80 mph....mph = miles per hour. neither train is going 80 miles per hour. one of the trains does go 80 *miles*, however

OpenStudy (anonymous):

the second train will go 75 miles.

OpenStudy (whpalmer4):

right! So how far apart are they at that point?

OpenStudy (anonymous):

sorry i meant 80 miles.

OpenStudy (whpalmer4):

another way to think of this problem is as two lines on a graph. The first train's position can be thought of as \(d_1= 40t\) if t is in hours. the second train's position can be thought of as \(d_2 = 60(t-0.75)\) if t is in hours.

OpenStudy (anonymous):

now i figured it out. thank you.

OpenStudy (whpalmer4):

then you plug in t = 2 in each equation to find out how far they've traveled, and subtract to get the distance apart. another related problem would be to ask when the second train catches the first. the graphical representation here is very helpful — where the two lines cross is directly above the time value

OpenStudy (whpalmer4):

My last "word problem" tip is to always, always check your answer to make sure it actually works :-)